Im having problems with recursive sequences (An= 3an-1), my secuence is 5,15,45,135... and im trying to make sure how to get 15 with these formula
Found 2 solutions by stanbon, tinbar:
Answer by stanbon(75887)

Im having problems with recursive sequences A(n) = 3(an-1), my sequence is 5,15,45,135... and im trying to make sure how to get 15 with these formula
---
Each succeeding number is 3 times its preceding number.
---
1st: 5
2nd: 3*5 = 15
3rd: 3*15 = 45
etc.
Cheers,
Stan H.

Answer by tinbar(133)

Your formula is correct. You just need to mention the first term A1 = 5.
Then your formula says for the nth term in the sequence, An, you take the previous term A(n-1) and multiply it by 3.
So the first term is 5, then the second term is 5*3 = 15, then the third term is 15*3 = 45 and so on